OpenStudy (anonymous):

An airplane flew 3 hours with a 40 mph head wind. The return trip with a tail wind of the same speed took 2 hours. Find the speed of the plane in still air. 180 mph 200 mph 220 mph 240 mph

OpenStudy (anonymous):

200 mph.. The first flight, he is pushed back 120 miles.. second flight he gains 80 miles... he does this in one hour.

OpenStudy (anonymous):

r t = d Solve the following for r:\[3(r-40)=2(r+40)\]r = 200 mph